OpenStudy (yb1996):

Use differentials to estimate the volume of thin metal spherical shells of thickness dr. If the metal has density 20gm/cm^3, how large must the radius be so that it floats in water (1gm/cm^3)? (Recall Archimedes Principle: the buoyant force is equal to the weight of the water being displaced.)

OpenStudy (mathmate):

r=outer radius in cm, dr=thickness in cm \(\rho_m = density of metal=20 g/mL\) \(\rho_w = density of water=1 g/mL\) Approximate volume of metal Vm=A*dr =\(4\pi~r^2dr\) mL Approximate mass of metal Mm=\(\rho_m ~Vm\) =\(20*4\pi~r^2~dr\) g Volume of sphere Vs=\((4/3)\pi~r^3\) mL If sphere is submerged, mass of water displaced \(M_w\)=volume of sphere * mass density of water = \((4/3)\pi~r^3 *\rho_w\) g For sphere to float, \(M_mg\le M_wg \) \(20*981*4\pi~r^2~dr\times 981\le (4/3)\pi~r^3 *1\times 981\) Solve for r in terms of dr.

OpenStudy (yb1996):

Wow, thank you so much for your help! I really appreciate the amount of work you put into that answer!

OpenStudy (yb1996):

I do have a question though, where did you get the number 981 from?

OpenStudy (mathmate):

981 is the acceleration due to gravity in cm/s^2. In this case, both densities are mass densities, so i need to multiply by g to get the weight (in force units) to satisfy Archimedes Principle. However, "g" cancels out on both sides, so the value does not really mater! :)

OpenStudy (mathmate):

*matter

OpenStudy (yb1996):

you multiplied 981 twice on the left side, but only once on the right side, so how would they cancel out?

OpenStudy (mathmate):

Oh, sorry, my bad, I wasn't paying attention, thought I missed it on both sides. There should be one on each side! My apologies! :(

OpenStudy (yb1996):

No problem, thanks again for being so helpful!

OpenStudy (mathmate):

mass * g = weight

OpenStudy (mathmate):

no problem! :)